OpenStudy (anonymous):

we have a vector sum on R^2 and define it as: (x1,y1)+(x2,y2)=(x1+x2,y1,y2). Explain why these scalar products aren't a good choice to make R^2 a vectorspace

OpenStudy (anonymous):

a) a(x,y)=(x,y) b) a(x,y)=(ax,0)

OpenStudy (anonymous):

my friend has this as his answer: \[1.v \neq v\]

OpenStudy (anonymous):

and \[1.(u_1,u_2)=(u_1,0) \neq (u_1,u_2)\]

OpenStudy (anonymous):

how did he get this?

OpenStudy (phi):

is this a typo (x1,y1)+(x2,y2)=(x1+x2,y1,y2) ? as written, your result is not in R^2

OpenStudy (anonymous):

odd..not a typo...

OpenStudy (anonymous):

i'm guessing he meant: (x1,y1)+(x2,y2)=(x1+x2,y1+y2)

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

please someone explain this..

OpenStudy (amistre64):

what do a and b mean? are they just options to consider?

OpenStudy (amistre64):

a(x,y) = (x,y) for some any a in R does not provide for an inverse element of (x,y)

OpenStudy (amistre64):

and the other doesnt appear to provide for an identity

OpenStudy (amistre64):

other than that ... thats all i can make of it

OpenStudy (anonymous):

i think its just question a) and b)

OpenStudy (amistre64):

then what i wrote prolly pertains .... its not that condusive to define the scalar multiplication in that manner since it would mess with identity and inverse porperties of vector spaces

OpenStudy (anonymous):

the answers my friend wrote, how did he come up with it? and what does he even mean with v?

OpenStudy (amistre64):

if you are asking questions about what your friend came up with ... you would prolly get better results by asking your friend.

OpenStudy (anonymous):

true, but i can only contact him next week

OpenStudy (amistre64):

his first result is off tho; 0(a,b) not= (0,0)

OpenStudy (anonymous):

alright..